# SpatialDreamer: Self-supervised Stereo Video Synthesis from Monocular Input

### v1

@@ 4.3.1 Stereo Image Synthesis
We visually compared the outputs of the different methods by generating a new view image given the target viewpoint. We qualitatively compare the newly synthesized images in Figure 5. 3D-photography fails to resolve the image distortion problem, and the image appears to be enhanced only slightly by the rendering. SynSin generates the most blurry details and loses a lot of information, compared with the input view. Neither AdaMPI nor SinMPI can generate fine enough details, and the results contain noticeable artifacts. CoPoNeRF tends to produce discrepancies from the original image in terms of fine-grained details, and it struggles with effectively handling scenes involving motion, as well as MVSplat. In contrast, the proposed method shows a clean and detailed output, especially in the area of depth information discontinuity.
@@
3D-photography [54$\dagger$ 0.641 0.446 11.08 9.533 0.116 0.386
SynSin [67$\dagger$ 0.740 0.645 21.22 19.55 0.087 0.104
@@
Table 1: Quantitative comparison on RealEstate10K dataset. $\dagger$ means re-implementing in our setting for fair comparison. The best results are in bold.
4.3.2 Stereo Video Synthesis
Moreover, we qualitatively compared the results of the stereo video synthesis with other methods. In addition to the quality issues of single-frame generation, the most important aspect in videos is the temporal consistency. We seek to generate videos that are free of flicker. The visualization of generated videos from different methods are illustrated in Figure 6, for each video, we show the scanline (the yellow line in figure) slice through the spatial-temporal volume. Our method achieve the smoothest trajectories, while other methods exhibit significant flickering. The most distinct regions are zoomed in and highlighted under the images to illustrate the details. In the first case, our method maintain the texture of the trees behind the woman. And the second case, only our method attain the smooth trajectories of the background (maintaining the straight line). And the last case, only our method maintain the texture of the black vertical lines. More results of generated frames are also presented in supplementary material.
@@ 4.3.2 Stereo Video Synthesis
4.4 Quantitative Results
Table 1 lists the SSIM, PSNR, and LPIPS scores for the newly synthesized images. The proposed method outperforms 3D-photography and SynSin in the major metrics by a wide margin. The method of Wang et al., CoPoNeRF and P-NVS perform slightly worse in terms of all metrics compared to the proposed method. Since the proposed method merges the spatial features and fills the occluded regions, it achieves the best LPIPS scores and shows competitive results in PSNR and SSIM when compared with AdaMPI and MVSplat. Overall, the results exhibit superior visual effects, which is consistent with the performance observed in the qualitative experiments. We further demonstrate the effectiveness of our method on video synthesis, the results are shown in Table 2. The proposed method achieves the best performance on both FVD and $E_{warp}$, which means our generated videos attain the best quality and temporal consistency.
@@ 4.5.1 Effectiveness of each components
We show the results in Table 3 on the RealEstate10K (stereo image) and collected stereo video datasets. CN means substituting our RefinerNet (RN) to ControlNet, SDS means the proposed stereo deviation strength. For image-level comparison, the results show that utilizing U-Net architecture only is not sufficient for stereo image synthesis and our RefinerNet is a better choice than ControlNet. Adding SDS further improves the geometric consistency. As visualized in Fig 7, solely relying on U-Net features results in distortion in the edge area and struggles with generating content properly. It also fails to preserve textural details near a plane boundary. Adding ControlNet often results in content inconsistency, presumably due to the failure of the paired view feature integration. RefinerNet handles this problem, showing it is a better architecture for stereo image generation. SDS further improves the consistency with the input and contain a more detailed image structure. As for video-level comparison, the quantitative results are shown in Table 3, adding TIL and DVG both attain the performance gain, and our final model achieves the best performance.
@@ 4.5.2 Data improvement by DVG
We show the generated stereo occlusion mask with inter-frame motion refinement in Fig 8. It can be observed that the occlusions appear more complete spatially and exhibit smoother in their temporal distribution. By employing DVG, we can acquire sufficient data to support the generation of spatio-temporally consistent videos.
@@ 4.5.3 Exploration of different depth estimation methods
We further compare different design choices of depth map including DepthAnything [71], Marigold [30], MiDaS [47] and ZoeDepth [5]. The results are available in supplementary materials. The proposed method can produce new viewpoint images that are realistic, seamless, and rich in detail, regardless of the depth estimation method used. This capability demonstrates the versatility and robustness of the proposed method, ensuring superior performance in various applications and environments.

### v2

@@ 4.3.1 Stereo Image Synthesis
We visually compared different methods for generating new view images from a target viewpoint. We qualitatively compared the newly synthesized images in Figure 5. 3D-photography fails to resolve the image distortion problem, and the image appears to be enhanced only slightly by the rendering. SynSin generates the most blurry details and loses a lot of information, compared with the input view. Neither AdaMPI nor SinMPI can generate fine enough details, and the results contain noticeable artifacts. CoPoNeRF tends to produce discrepancies from the original image in terms of fine-grained details, and it struggles with effectively handling scenes involving motion, as well as MVSplat. In contrast, the proposed method shows a clean and detailed output, especially in the area of depth information discontinuity.
4.3.2 Stereo Video Synthesis
Moreover, we qualitatively compared the results of the stereo video synthesis with other methods. In addition to the quality issues of single-frame generation, the most important aspect in videos is the temporal consistency. We seek to generate videos that are free of flicker. The visualization of generated videos from different methods are illustrated in Figure 6, for each video, we show the scanline (the yellow line in figure) slice through the spatial-temporal volume. Our method achieve the smoothest trajectories, while other methods exhibit significant flickering. The most distinct regions are zoomed in and highlighted under the images to illustrate the details. In the first case, our method attains the smooth trajectories of the background (maintaining the straight line) . And the second case, only our method preserves the texture of the trees behind the woman. And the last case, only our method preserves the texture of the black vertical lines. More results of generated frames are also presented in supplementary material.
4.4 Quantitative Results
Table 1 lists the SSIM, PSNR, and LPIPS scores for the newly synthesized images. The proposed method outperforms 3D-photography and SynSin in the major metrics by a wide margin. The method of Wang et al., P-NVS and NVSVDE-Net perform slightly worse in terms of all metrics compared to the proposed method. Since the proposed method merges the spatial features and fills the occluded regions, it achieves the best LPIPS scores and shows competitive results in PSNR and SSIM when compared with AdaMPI and CoPoNeRF. MVSplat outperforms other methods when $t$=10 but is inferior to ours when $t$=5. Overall, the results exhibit superior visual effects, which is consistent with the performance observed in the qualitative experiments. We further demonstrate the effectiveness of our method on video synthesis, as shown in Table 2. The proposed method achieves the best performance on both FVD and $E_{warp}$, which means that our generated videos attain the best quality and temporal consistency.
Table 1: Quantitative comparison on RealEstate10K dataset. $\dagger$ means re-implementing in our setting for fair comparison. When $t$=5, the current frame serves as the left view and the frame at time$+5$ serves as the right view; the same approach applies for $t$=10. The best results are in bold.
@@ 4.4 Quantitative Results
3D-photography [56$\dagger$ 0.641 0.446 11.08 9.533 0.116 0.386
SynSin [69$\dagger$ 0.740 0.645 21.22 19.55 0.087 0.104
@@ 4.5.1 Effectiveness of each components
We show the results in Table 3 on the RealEstate10K (stereo image) and collected stereo video datasets. CN means substituting our RefinerNet (RN) to ControlNet, SDS means the proposed stereo deviation strength. For image-level comparison, the results show that utilizing U-Net architecture only is not sufficient for stereo image synthesis and our RefinerNet is a better choice than ControlNet. Adding SDS further improves the geometric consistency. As visualized in Figure 7, solely relying on U-Net features results in distortion in the edge area and struggles with generating content properly. It also fails to preserve textural details near a plane boundary. Adding ControlNet often results in content inconsistency, presumably due to the failure of the paired view feature integration. RefinerNet handles this problem, showing it is a better architecture for stereo image generation. SDS further improves the consistency with the input and contain a more detailed image structure. As for video-level comparison, the quantitative results are shown in Table 3, adding TIL and DVG both attain the performance gain, and our final model achieves the best performance.
@@ 4.5.2 Data improvement by DVG
We show the generated stereo occlusion mask with inter-frame motion refinement in Figure 8. It can be observed that the occlusions appear more complete spatially and exhibit smoother in their temporal distribution. By employing DVG, we can acquire sufficient data to support the generation of spatio-temporally consistent videos. More analyses are available in the supplementary materials.
@@ 4.5.3 Exploration of different depth estimation methods
We further compared different design choices of depth map including DepthAnything [73], Marigold [31], MiDaS [49], ZoeDepth [5] and DepthCrafter [26]. The results are available in supplementary materials. The proposed method can produce new viewpoint images that are realistic, seamless, and rich in detail, regardless of the depth estimation method used. This capability demonstrates the versatility and robustness of the proposed method, ensuring superior performance in various applications and environments.
4.6 Limitations
SpatialDreamer effectively creates stereo video from a single input but faces limitations. It relies on depth maps for rendering, affecting stereoscopic quality due to depth accuracy and distribution. Future research could focus on integrating implicit depth to address this [36]. Additionally, the model’s large parameter size limits real-time performance.
@@ 5 Conclusion
Supplementary Material
6 Image Synthesis
6.1 Visual Comparison with Other Methods
More results of comparison with other methods are shown in Figure 9.
Figure 9: Visual result on the RealEstate10K dataset.
6.2 Visual Result of Ablation Study
More results of ablation study are shown in Figure 10. By zooming in on the images, one can more clearly discern the differences in detail, including artifacts, noise, and discontinuities.
Figure 10: Visual result of ablation study.
6.3 Visual Comparison of Different Depth Estimation Methods
As shown in Figure 11, DepthAnything is effective in capturing fine details and maintaining the consistency of depth across different scenes, while Marigold produces sharp and detailed depth maps. MidaS produces smooth and coherent depth maps, and ZoeDepth shows a balanced approach to detail preservation and depth accuracy, especially better reconstruction of flat surfaces. The proposed method consistently delivers excellent results in generating new viewpoint images, regardless of the depth estimation method used.
Figure 11: Visual comparison of different depth estimation methods. The zoomed image represents occluded regions.
7 Video Synthesis
7.1 Computational Cost
As shown in Table 4 with our 2.485 billion model.
Table 4: Cost for a 1024x1024, 30-frame video on an A800 GPU.
3D-photography Webui-depthmap P-NVS CoPoNeRF NVS-Solver MVSplat Proposed
DVG - - - - - - 51s
Inference - - - - - - 1059s
Total 10800s 488s 1503s 519s 833s 41s 1110s
7.2 Quantitative Impact of Depth and Motion Estimation
Table 5 and Table 6 show the impact of motion estimation methods and depth estimation methods, respectively.
7.3 Quantitative comparison
Table 7 shows that our DVG/TIL module remains valid even with video depths.
Table 5: Quantitative comparison of motion estimation methods.
PWC-Net RAFT(Proposed) SEA-RAFT
[…]
$E^{*}_{warp}$downarrow 3.661 3.374 3.302
Table 6: Quantitative comparison of depth estimation methods.
DepthAnything Marigold ZoeDepth DepthCrafter MiDaS(Proposed)
[…]
$E^{*}_{warp}$downarrow 3.390 3.572 3.365 3.333 3.374
Table 7: Ablation study of the DepthCrafter method
[…]
127.6 3.593
76.27 3.393
81.91 3.341
62.50 3.333
7.4 Comparison with Other Methods
Figure 12: Quantitative comparison with the other methods.The first three columns are adjacent frames with t, t+1, t+2, and the last column is the non-adjacent frame. The proposed method not only generates accurate and consistent content in the occluded regions among the adjacent frames, but also maintains consistency with the visible parts in the non-adjacent frame, as highlighted in the orange boxes.
Video results compared with other methods are provided in folder “$\mathbb{compare\_to\_others}$” and the corresponding visual comparison is shown in Figure 12. The meaning of the file name is explained below:
  • input.mp4: input video
  • 3d-photo.mp4: video result using 3D-photography method
  • webui.mp4: video result of webui-depthmap method
  • P-NVS.mp4: video result of P-NVS method
  • coponerf.mp4: video result of CoPoNeRF method
  • nvs-solver.mp4: video result of NVS-Solver method
  • AVP.mp4: video result of Apple Vision Pro
  • ours.mp4: video result of our method
7.5 Videos Under the Target Viewpoint
The newly synthesized video results of our method are provided in folder “$\mathbb{target\_view\_video}$”.
7.6 Stereo Video Results
Furthermore, the final side-by-side video results are provided in folder “$\mathbb{stereo\_video}$”, which can be watched on VR device(AVP, Quest, Pico, etc.).